Suppose that the number of drivers who travel between a particular origin and destination during a designated time period hasPlease explain step by step
Answer #1

x~possion(\mu=20)

P(X = ?) - e-* Hl

Ans a:

P(X<11)

107 * 07- P(X <11) 0=

P(X <11) = Σ e-20 και 20 -= 0.021 ! =0

Ansb

P(X>26)

P(X>26)=1-P(x<=26)

26 P(X < 26) e-20 * 20 -= 0.922

P(X>26)=1-0.922

=0.078

Ans c:

26 P(11 < X < 26) => e-20 * 201 -= 0.913 I=11

#Probability of strictly between 11 and 26

= 25 -20 $ 20 P(11 < X < 26) = – = 0.866 . T=12

Ans d:

here mean =20

and std deviation =sqrt(20)=4.47

hence P(within 2 std deviation from mean )=P(20-2*4.47<X<20+2*4.47)=P(11.06 <X<28.94)=Σ - 2020 /r1 T=12

=0.944